# Generalizable LLM Learning of Graph Synthetic Data with Reinforcement Learning

### v2

@@ E.1 Additional Translation Between Single-Step and Multi-Step
We deliberately probe the model with a single constraint per prompt for our Proscript single-step setting. Results are shown in Table 6. When using a significance threshold of $\alpha=0.05$, we find that the performance difference between the two experimental groups is all statistically significant, with 5 out of 6 analysis results showing that multi-step is doing much better. On one hand, SFT using synthetic data leads to decreased multi-step reasoning capabilities on real-world tasks; on the other hand, the model’s performance increases, with the help of real-world SFT or synthetic RL, but comes with a more severe hallucination on multi-step performance, as the performance gap between single-step and multi-step is significant. We can cautiously conclude that in Proscript, performance gains rise from more severe hallucination after real-world SFT or synthetic RL, rather than increased graph reasoning capabilities. Proscript’s results align with previous results of StrategyQA and Knowledge Crosswords, and further emphasize the importance of understanding and investigating hallucination-free and explainable intermediate steps in multi-step reasoning tasks.
Qwen2.5-7B-Instruct Single-step Multi-step Proportion z-test p-val Zero-Shot 0.575 0.592 0.046 Synthetic SFT 0.627 0.488 7.35E-61 Best RL: GRPO-P w/o SFT 0.573 0.605 1.00E-4 Real-world SFT 6k 0.681 0.851 1.24E-124 Real-world SFT 1.2k 0.613 0.795 1.51E-123 Real-world SFT 0.3k 0.607 0.729 2.01E-52
@@ E.2 Compositionality Gap Implementation Details
Figure 4: Performances on Blocksworld Verification task grouped by the length of action steps needs to be verified. With synthetic SFT and real-world SFT, the performance on verifying plans with longer steps gradually increases, while using RL alone cannot produce the result.
E.3 Synthetic SFT’s Role on Real-world Tasks
@@
Results in Figure 4 show surprising results: with the help of synthetic SFT and real-world SFT, the model’s performance on long sequence plans increases, while only using RL methods cannot provide this result. We believe that this result comes from the help of strong supervision using strict format SFT data without any hallucination errors. This result also strengthens the idea that providing correct reasoning steps for the correct tasks should generally be helpful, while the best method for supervision still require future research.

### v3

@@ E.1 Additional Translation Between Single-Step and Multi-Step
We deliberately probe the model with a single constraint per prompt for our Proscript single-step setting. Results are shown in Table 6. When using a significance threshold of $\alpha=0.05$, we find that the performance difference between the two experimental groups is all statistically significant, with 5 out of 6 analysis results showing that multi-step is doing much better. On one hand, SFT using synthetic data leads to decreased multi-step reasoning capabilities on real-world tasks; on the other hand, the model’s performance increases, with the help of real-world SFT or synthetic alignment, but comes with a more severe hallucination on multi-step performance, as the performance gap between single-step and multi-step is significant. We can cautiously conclude that in Proscript, performance gains rise from more severe hallucination after real-world SFT or synthetic alignment, rather than increased graph reasoning capabilities. Proscript’s results align with previous results of StrategyQA and Knowledge Crosswords, and further emphasize the importance of understanding and investigating hallucination-free and explainable intermediate steps in multi-step reasoning tasks.
Qwen2.5-7B-Instruct Single-step Multi-step Proportion z-test p-val Zero-Shot 0.575 0.592 0.046 Synthetic SFT 0.627 0.488 7.35E-61 Best Alignment: GRPO-P w/o SFT 0.573 0.605 1.00E-4 Real-world SFT 6k 0.681 0.851 1.24E-124 Real-world SFT 1.2k 0.613 0.795 1.51E-123 Real-world SFT 0.3k 0.607 0.729 2.01E-52
@@ E.2 Compositionality Gap Implementation Details
Figure 4: Performances on Blocksworld Verification task grouped by the length of action steps needs to be verified. With synthetic SFT and real-world SFT, the performance on verifying plans with longer steps gradually increases, while using alignment alone cannot produce the result.
E.3 Correlation Analysis
We investigate the correlation between various performance and reward metrics. First, to better understand the reward design’s effect, we analyze the correlation between the synthetic task’s accuracy performance and the reward score achieved by the model. The Pearson correlation coefficient is 0.951: our reward design is highly correlated with the performance of the model, proving the reward function’s effectiveness. Second, we analyze the correlation between synthetic performance and real-world performance (denoted by the mean performance of all real-world tasks). The correlation is -0.336, which aligns with our finding that performance gains provided by synthetic training are mixed across models and different tasks. Detailed results are shown in Appendix B Figure 3.
E.4 Synthetic SFT’s Role on Synthetic Tasks
SFT stage acts as different roles for different alignment methods for synthetic tasks, as shown in Table 1. For on-policy method GRPO, it may decrease the robustness of the trained model as 2 out 4 settings (SFT + GRPO) cannot achieve comparable results compared to those GRPO trained models trained without SFT stage, leading to limitation of the model’s performance if SFT stage is enforced before GRPO. However, for off-policy method DPO, SFT can somehow bring advantage to the model, with 2 out of 4 settings (SFT + DPO) achieved better results compared to only those only trained using SFT, while those models directly trained using DPO cannot achieve better results than SFT models.
E.5 Synthetic SFT’s Role on Real-world Tasks
@@
Results in Figure 4 show surprising results: with the help of synthetic SFT and real-world SFT, the model’s performance on long sequence plans increases, while only using alignment methods cannot provide this result. We believe that this result comes from the help of strong supervision using strict format SFT data without any hallucination errors. This result also strengthens the idea that providing correct reasoning steps for the correct tasks should generally be helpful, while the best method for supervision still require future research.